OpenStudy (anonymous):

What is the equation of the line that passes through the points (−1, 3) and (1, 11)? 4x − y = 1 4x + y = 1 x − 4y = −13 4x − y = −7

OpenStudy (anonymous):

4 is correct. \[y = 4x + 7\]

OpenStudy (aroub):

y=mx+b , m= 4 11=4(1)+b 11=4+b 7=b y=4x+7 4x − y = −7